### v2

# DocReward: A Document Reward Model for Structuring and Stylizing Junpeng Liu1 Yuzhong Zhao211footnotemark: 1Bowen Cao1Jiayu Ding3Yilin Jia4Tengchao Lv5 Yupan Huang5Shaohan Huang5Nan Yang5Li Dong5Lei Cui5Tao Ge5Xun Wang5 Huitian Jiao5Sun Mao5FNU Kartik5Si-Qing Chen5Wai Lam1Furu Wei5 1CUHK

Abstract
Recent advances in agentic workflows have enabled the automation of tasks such as professional document generation.
However, they primarily focus on textual quality, neglecting visual structure and style, which are crucial for readability and engagement.
This gap stems mainly from a lack of effective reward models capable of guiding agents toward producing documents with high structural and stylistic professionalism.
To address this, we propose DocReward, a Document Reward Model that evaluates documents based on their structure and style.
The model is trained under a textual-quality-agnostic framework to assess professionalism without being influenced by textual quality. To achieve this,
we construct a multi-domain dataset DocPair of 117K paired documents, covering 32 domains and 267 document types, each comprising a high- and low-professionalism document with identical content but different structure and style.
This setup enables the model to evaluate professionalism comprehensively and independently of textual quality.
DocReward is trained using the Bradley-Terry loss to score documents, penalizing predictions that contradict the annotated ranking.
On a manually annotated benchmark, DocReward outperforms GPT-5 by 14.6 percentage points in accuracy.
Extrinsic RL experiments further validate its effectiveness in guiding professional document generation.
DocReward: A Document Reward Model for Structuring and Stylizing
Junpeng Liu1††thanks: Equal contribution. Work done during internship at Microsoft Research.
Yuzhong Zhao211footnotemark: 1
Bowen Cao1
Jiayu Ding3
Yilin Jia4
Tengchao Lv5Yupan Huang5Shaohan Huang5Nan Yang5Li Dong5Lei Cui5Tao Ge5Xun Wang5Huitian Jiao5Sun Mao5FNU Kartik5Si-Qing Chen5Wai Lam1Furu Wei51CUHK
2UCAS
3XJTU
4UMich
5Microsofthttps://aka.ms/GeneralAI
1 Introduction
Figure 1: DocReward automatically assesses document professionalism according to their structure and style, assisting existing agentic workflows for more professional document generation (left). It outperforms GPT-5 by 14.6% in human preference accuracy (right).
Recent advances in agentic workflows have automated various complex tasks, such as code generation (Peng et al., 2023; Cherny and Anthropic, 2025; Hong et al., 2024), image generation (comfyanonymous, 2025), visual understanding (Zheng et al., 2025; Marsili et al., 2025), math reasoning (Yan et al., 2025), and travel planning (Xie et al., 2024). A key focus of agentic workflows is the generation of professional documents, including works like deep research (OpenAI, 2025b; Liang et al., 2025; Qwen, 2025) and technical documentation generation (Dvivedi et al., 2024). However, existing research on professional document generation primarily focuses on improving textual quality, often neglecting the visual structure and style—both of which are crucial for document professionalism.
A well-organized structure facilitates seamless navigation for readers, while a consistent style makes the content more readable and engaging. Together, these aspects help convey information more clearly and effectively.
The neglect of structure and style mainly stems from the lack of effective reward models, which are capable of guiding agentic workflows toward producing documents with professional structure and style.
However, building a reward model capable of providing a robust evaluation of visual structure and style is non-trivial, as it requires both comprehensiveness and textual-quality-agnosticism. Specifically, comprehensiveness refers to the ability to evaluate documents across diverse types, qualities, structures, and styles, while textual-quality-agnosticism, in this context, means that the model does not evaluate the inherent quality of the textual content itself, but instead assesses how well the structure and style of a document stand out, given the fixed content.
To achieve this, we propose DocReward, a Document Reward Model specialized in assessing document professionalism in structure and style, as illustrated in Figure 1.
The model is trained under a textual-quality-agnostic framework. Specifically,
we construct DocPair, a multi-domain dataset of 117K document pairs across 32 domains and 267 types. Each pair comprises a high-professionalism sample and its low-professionalism counterpart; these documents share identical content but differ in structure and style.
The construction of DocPair involves three phases:
1) Curating High-Quality Professional Documents. We curate a collection of high-quality documents characterized by professional structure and style from diverse domains, such as government, education, and science.
2) Expanding Source Documents via Agents. Next, we extract the textual content and employ multiple generation agents to re-produce documents that preserve the original textual content while exploring diverse structural and stylistic variations.
3) Ranking Documents. The ranking labels are determined by a combination of human-verified heuristics and an oracle-based annotation method. This provides a scalable way to capture elements of structural and stylistic professionalism.
Based on the constructed dataset, we train DocReward to take rendered document pages as input and output a score reflecting the document’s professionalism in structure and style.
The predicted scores of paired documents are optimized using the Bradley-Terry loss (Bradley and Terry, 1952; Ouyang et al., 2022), which penalizes violations of the annotated ranking order.
To demonstrate the superiority and utility of DocReward, we conduct both intrinsic and extrinsic evaluations.
For the intrinsic evaluation, we establish a benchmark DocPairBench of 1443 human-annotated pairs across multiple domains. Human annotators ranked each pair based on the professionalism of the documents’ structure and style. Notably, as shown in Figure 1 (right), DocReward outperforms GPT-4o (Hurst et al., 2024) and GPT-5 (OpenAI, 2025a) by 27.4 and 14.6 percentage points, respectively, in accuracy on DocPairBench, demonstrating its superiority over existing baselines.
For the extrinsic evaluation, we evaluate DocReward through two complementary experiments.
1) Best-of-N. DocReward is used as a re-ranking model for improving agentic workflow without changing the agent itself. Human evaluation reveals that DocReward as a reward model achieves a significantly higher win rate of 60.8%, compared to GPT-5’s 37.7%.
2) Reinforcement Learning. We further demonstrate the utility of DocReward as the reward model for reinforcement-learning of both open- and closed-source models. This integration improves the document generation performance of Qwen2.5-Coder (Hui et al., 2024) and GPT-4o regarding structure and style.
Our main contributions are as follows:
We propose a textual-quality-agnostic framework for document reward modeling. By decoupling structure and style from textual content, it enables the reward model DocReward to accurately capture complex structural and stylistic elements.
DocPair: A large-scale multi-domain dataset comprising 117K document pairs across 32 domains and 267 document types, designed to equip DocReward with comprehensiveness and textual-quality-agnosticism.
Comprehensive experiments demonstrate that DocReward not only outperforms GPT-5 in assessing structure and style, but also serves as an effective reward model in reinforcement learning for agentic workflows.
2 Textual-Quality-Agnostic Framework
Document professionalism is characterized by its textual content, structure, and style. Although large language models excel at evaluating textual quality, they are limited in assessing structure and style.
To bridge this gap, we propose a textual-quality-agnostic framework that trains reward models to evaluate professionalism independently of textual quality.
In this section, we formulate the framework and define its objectives.
Let denote a set of documents, where each document consists of textual content and rendered images . The document reward model assigns scores to documents that share the same textual content, such that these scores reflect their structural and stylistic quality. This process is formalized as follows:
where “” is a predefined similarity function that measures the agreement between the true and predicted quality orders. “” returns the indices of documents sorted by their predicted scores.
denotes the true indices reflecting the relative ranking of the documents in terms of structure and style.
Essentially, the term “” ensures that the model evaluates professionalism in a textual-quality-agnostic manner, as it processes identical textual content from paired documents.
In this paper, document professionalism in structure and style is defined as follows:
1) Structure: Proper use of white space, appropriate margins, clear section breaks, well-structured text alignment, adequate paragraph spacing, proper indentation, inclusion of page headers and footers, and logical, coherent organization of content.
2) Style: Appropriate font choices (type, size, color, readability), clear heading styles, effective use of emphasis (bold, italics), bullet points, numbering, and consistent formatting.
By optimizing based on these factors, we obtain a reward model capable of assessing structural and stylistic professionalism in a comprehensive and textual-quality-agnostic way.
3 DocReward
Based on the proposed textual-quality-agnostic framework, we train DocReward, a reward model specializing in assessing the structural and stylistic professionalism of documents. DocReward is trained on DocPair, a diverse dataset of 117K document pairs (Section 3.1), and is optimized with a preference-based objective for structural and stylistic assessment (Section 3.2). The following sections detail the data construction pipeline and model design.
Figure 2: The data construction pipeline for DocReward.
3.1 Data Construction
As shown in Figure 2, we first collect a set of high-quality real-world source documents.
The source documents are then expanded by multiple generation agents, and the resulting documents are grouped by shared textual content.
Finally, each group of documents is annotated with a ranking in terms of structure and style quality.
The construction procedure is detailed step by step below:
Curating High-Quality Professional Documents. As illustrated in Figure 2 (top), we first curate a corpus of human-authored Microsoft Word documents that spans both highly formal institutional writing and everyday professional communication. We draw on two complementary sources:
1) Government corpora: GovDocs1 (Garfinkel et al., 2009) and NapierOne (Davies et al., 2022) are authoritative document collections sourced from government and public institutions, covering reports, forms, guidelines, and other professional materials with consistent structure and style.
2) Web document corpus: CommonCrawl111https://commoncrawl.org/, covering business, education, healthcare, etc, enriching structural and stylistic diversity.
To ensure high quality, we apply a preprocessing and filtering pipeline before data construction. First, all files are converted to DOCX format to enable programmatic access and modification via python-docx222https://python-docx.readthedocs.io/. Next, we discard extreme or malformed cases (exceeding 20 pages, files larger than 1 MB dominated by images, and files smaller than 10 KB with trivial content). To efficiently reduce residual noise, we employ GPT-5 as a rigorous automated heuristic to flag poor structure/style on a scale; documents scoring above 8 are retained. A manual inspection of 200 randomly sampled retained documents confirms that this automated filter preserves high-quality professional samples.
The domain distributions of the filtered corpus are shown in Figure 3. The corpus spans 32 domains and over 267 document types, demonstrating substantial breadth and diversity, and it provides a high-quality foundation for constructing subsequent document pairs.
Figure 3: Top 10 Document Domain Distribution (Total: 32).
Expanding Source Documents via Agents.
As shown in Figure 2(middle), to obtain documents with the same textual content but different structure and style, we construct two types of agents to synthesize documents given the textual content (and rendered pages) of the source documents. To further increase the diversity of the synthesized documents, each agent can be empowered by different LLMs. The two agents are detailed as follows:
1) Textual Content to Document.
The textual content is first extracted from the source documents, discarding all formatting, styling, and layout information. Then, advanced generation models (, GPT-4o, OpenAI o1 (OpenAI, 2024), Claude Sonnet 4 (Anthropic, 2025), and GPT-5) are used to synthesize DOCX documents via python-docx.
2) Refinement for Better Structure and Style.
To further improve the structure and style of synthesized documents, we refine them by comparing them with the original human-authored documents in terms of structure and style. The refinement process consists of two stages:
1) Agents are provided with the python-docx code, rendered pages, and structured textual representation of the synthesized document, along with the rendered pages of the original human-authored document, to generate a refinement plan.
2) Based on the plan, the agents modify the python-docx code to produce refined documents with improved structure and style.
The synthesized documents are grouped with their originals to facilitate subsequent processing. Please refer to Appendix A.3 and A.6 for details.
Ranking Documents.
As shown in Figure 2 (bottom), the collected documents within each group share identical textual content and are organized into pairs. The annotation task is to assess the relative professionalism in terms of structure and style for each pair, which is carried out under the following two cases:
1) Real vs. Synth.
When comparing an original document with its agent-generated counterparts, the human-authored version is designated as the winner. This heuristic is grounded in our preliminary human inspection, which indicates that current state-of-the-art models (e.g., GPT-5, Claude Sonnet 4) produce documents with structure and stylistic quality inferior to high-quality human-authored documents curated by the rigorous filtering pipeline (described in 3.1).
2) Synth. vs. Synth. For pairs where both documents are agent-generated, manual annotation is not scalable. We therefore employ a closed-source model as a proxy for human judgment. To mitigate its intrinsic biases, we adopt an oracle setting: the model is provided with a document triplet , using the human-authored as a reference anchor. This transforms the annotation from a subjective preference task into an objective similarity matching against the ground-truth document (see subsection A.6).
We validated this approach on 120 pairs, where it demonstrated 92.5% alignment with human experts, confirming its effectiveness.
Note that the above strategy is cost-effective and necessary for large-scale training data construction, while the test split, DocPairBench, is constructed through strict human annotation. The data statistics of DocPair are shown in Table 1.
Domains
Types
Doc.
Avg. Page
Pairs
32
267
69,137
3.2
117,108
Table 1: Data statistics of the constructed DocPair.
3.2 Model Structure and Optimization
We use Qwen-2.5-VL (Bai et al., 2025) as the base model, which takes a multi-page document as input and outputs a scalar score via an added regression head (details provided in Appendix A.2).
Given a preference pair consisting of a winner and a loser , the reward model assigns scores to both documents and is optimized using the Bradley–Terry (BT) loss:
where . This objective encourages higher scores for preferred documents.
4 Experiments
We first present DocPairBench (Section 4.1) and then conduct intrinsic and extrinsic evaluations as shown in Figure 4 (Section 4.2 and 4.3).
Figure 4: Three evaluation settings for DocReward.
4.1 DocPairBench
A subset of the curated documents in Section 3.1 is set aside as evaluation documents. To diversify the evaluation samples, we consider the following six types of documents using the method described in Section 3. Four of them are obtained via the Textual Content to Document agent, which generates DOCX documents using different LLMs (, GPT-4o, OpenAI o1, Claude Sonnet 4, and GPT-5). One type comes from the Refinement for Better Structure and Style agent, where GPT-5 is employed to refine synthesized documents. The last type consists of the curated human-authored documents. Together, these six types constitute the origins of samples in our benchmark.
For each set of documents sharing the same content but differing in structure and style, human experts meticulously rank their quality based on structure and style. To facilitate model evaluation, these ranked relationships are converted into a total of 1443 comparison pairs, each consisting of two documents and a binary label indicating the preferred one. To ensure the quality of human annotation, we evaluate annotation consistency among human annotators using Cohen’s Kappa and observe a high agreement of 83.4.
The details of annotation protocol and reliability validation are presented in Appendix A.4.
Table 2: Accuracy of Models on the DocPairBench benchmark across ten domains: Government (Gov.), Education (Edu.), Non-Profit Organization (NPO), Medical (Med.), Science (Sci.), Legal (Leg.), Business (Bus.), Academic (Acad.), Technology (Tech.), and Other (Oth.).
As presented in Table 2, we evaluate all methods under two settings, namely Pairwise and Pointwise. In the pairwise setting, models are given a document pair and asked to select the one with better structure and style, while in the pointwise setting, documents are scored independently. Performance is measured by accuracy with respect to human annotations. On DocPairBench, DocReward-3B and DocReward-7B achieve substantial improvements over strong baselines including Qwen2.5-VL, GPT-4o, Claude Sonnet 4, and GPT-5. In particular, DocReward-7B attains an overall human preference accuracy of 82.3 percent, surpassing the strongest closed source model (GPT-5) by 14.6 points. These results indicate that DocReward captures structural and stylistic quality signals that existing language models often overlook.
4.3 Improving Document Generation with DocReward
To demonstrate the utility of DocReward, we conduct two complementary experiments.
Rewards
Win
Lose
Tie
Random
24.6
66.2
9.2
GPT-5
37.7
40.0
22.3
DocReward
60.8
16.9
22.3
Table 3: Best-of-N evaluation results. DocReward shows utility for professional document generation.
Reward Types
Success
ROUGE-L
DocReward
Rank
Qwen2.5-Coder
30.0
20.61
0.0663
4.58
+ rule
98.0
97.94
0.1785
4.06
+ DocReward
100.0
97.95
0.3046
2.84
GPT-4o
52.0
48.73
0.2682
3.18
+ rule
66.0
62.15
0.3189
2.70
+ DocReward
78.0
74.33
0.4486
2.02
Table 4: Results of the reinforcement learning experiments. “DocReward” denotes the sigmoid-normalized DocReward score. “Rank” denotes the average ranking assigned by human annotators.
Best-of-N.
A document agent generates candidates from the same textual content, and a reward model selects the best one based on its score. We compare three reward models including “Random”, GPT-5 and DocReward with human annotators ranking the selected outputs by structure and style. As shown in Table 4, DocReward consistently outperforms the baselines, indicating that its reward better aligns with human preferences. This evaluation shows that integrating DocReward into a document agent improves output quality without modifying the agent itself. Details are provided in Appendix A.7.
Reinforcement Learning.
We aim to enhance document generation agentic workflows that take plain text as input and generate professional documents. We consider two kinds of rewards: 1) that penalizes documents that either result from invalid python-docx code or differ from the input text after execution. Specifically, if the code executes successfully, then ; else will be zero. 2) that penalizes documents with poor structure and style. Overall, the reward assigned to each generated document is defined as:
where is a hyperparameter to balance the rewards, is the operation to regularize the value range of DocReward to , and represents whether is larger than a threshold.
For open-source models, we adopt GRPO (Shao et al., 2024) as the reinforcement learning algorithm, while employing training-free GRPO (Cai et al., 2025) for closed-source models. After RL, human annotators are asked to rank the documents produced by six model variants. The evaluation criterion is the professionalism of the document’s structure and styling. The experimental results are shown in Table 4. Rule-based rewards substantially improve the success rate of document generation for both Qwen2.5-Coder and GPT-4o. Incorporating DocReward as reward further enhances the performance of both open- and closed-source models, leading to higher success rates, improved ROUGE-L scores, and better average rankings. These results demonstrate that DocReward serves as an effective reward model for professional structure and style. Figure 5 presents visualization of documents generated by different models.
Figure 5: Visualization of documents generated by various models. “Execution error” denotes instances where the generated code encountered runtime failures, preventing the final document from being rendered.
Table 6: Cross-lingual robustness evaluation. denotes the performance gap between English and non-English scenarios. DocReward demonstrates strong cross-lingual robustness.
Out-of-Domain Evaluation. Table 6 reports the out-of-domain results of different models. Firstly, DocReward-7B (77.5) remains superior to all baseline models, including the closed-source model GPT-5 (68.4). This trend is consistent with the in-domain results. Secondly, the performance of DocReward-7B decreases by merely 4.8 percentage points when transitioning from in-domain to out-of-domain evaluations. Such a small performance gap indicates that DocReward generalizes effectively to unseen domains.
Cross-Lingual Robustness: To evaluate robustness across languages, we conduct experiments in French, Spanish, and Russian, with results shown in Table 6. Firstly, in non-English scenarios, DocReward-7B model achieved a remarkable score of 77.9, substantially outperforming all baseline models. Secondly, all models, including the baselines, exhibited performance degradation in non-English settings. The performance drop of DocReward () is even smaller than that of closed-source models GPT-4o () and GPT-5 (), indicating that DocReward demonstrates strong cross-lingual robustness.
4.5 Visualization of Attention Map
To analyze DocReward’s decision-making process, we conduct probing experiments on attention maps. As shown in Figure 6, the model attends more to structural and formatting cues than semantic content when assessing document professionalism. Specifically, attention focuses on headings and numbering (Figure 6), page headers and footers (e.g., “CS-66”, “DEC. 2006”), bullet points (Figure 6), and table grids (Figure 6), reflecting sensitivity to structure clarity, formatting consistency, alignment, and readability. Additionally, attention to page corners suggests that uniform margins and balanced whitespace are important indicators of professional layout design.
Figure 6: Visualization of attention maps. DocReward captures structural and stylistic elements, such as headings, alignment, and whitespace, in its evaluation of document professionalism.
5 Related Work
Aesthetic and Professionalism Assessment.
In graphic design, AesthetiQ (Zhang et al., 2024) utilizes multimodal LLMs as preference evaluators to align layout generation with aesthetic requirements, while diffusion-based methods such as LACE (Li et al., 2023) introduce differentiable constraints to directly optimize layout attributes.
For web and mobile interfaces, systems like Calista (Yu et al., 2019) and Android UIs (Fu et al., 2024) use explicit ratings and pairwise comparisons to model visual appeal, showing correlations with usability.
Additionally, photo aesthetics are modeled using layout-aware CNNs such as A-Lamp (Li et al., 2018), and similar techniques extend to video (Liu and Yu, 2023).
These studies focus on images or UI interfaces rather than multi-page documents, where professionalism depends on both structure and style.
Document AI.
Document AI research mainly targets semantic parsing and content understanding.
Models such as LayoutLM (Xu et al., 2020) and ReLayout (Jiang et al., 2024), along with OCR-based pipelines (Subramani et al., 2020), identify logical elements such as headings, tables, and semantic groups to support information extraction and classification.
Recent work also explores automatic document or layout generation (Lin et al., 2023; Tang et al., 2023; Tian et al., 2025), but evaluation has primarily been limited to content correctness or basic formatting.
As a result, the assessment of document professionalism—particularly visual structure and style—remains largely unexplored.
Preference Learning and Reward Models.
A major challenge in professionalism assessment is acquiring feedback signals that reflect human judgment.
Preference-based reward modeling addresses this issue by training on pairwise comparisons to approximate preferences, forming the basis of alignment methods like RLHF (Stiennon et al., 2020) and DPO (Rafailov et al., 2023).
6 Conclusion
In this paper, we introduced DocReward, a document reward model designed to assess structural and stylistic professionalism.
To enable professionalism assessment without being influenced by textual quality, we propose a textual-quality-agnostic framework for document reward modeling. Then, a multi-domain preference dataset, DocPair, consisting of 117K paired documents, is constructed to enable the training of DocReward using the Bradley-Terry loss.
Comprehensive experiments demonstrate that DocReward not only outperforms GPT-5 in assessing structure and style, but also serves as an effective reward model in reinforcement learning for agentic workflows.
7 Limitations
A primary limitation is that DocReward is designed as a scalar reward model, providing only a numerical score to represent the overall structural and stylistic quality of a document. It currently cannot generate natural language feedback explaining why a document is perceived as unprofessional. Therefore, extending DocReward from a scalar model to an interpretable reward model that can produce both quantitative scores and qualitative diagnostic rationales remains a key focus of our future work.
Bai et al. (2025)
Shuai Bai, Keqin Chen, Xuejing Liu, Jialin Wang, Wenbin Ge, Sibo Song, Kai Dang, Peng Wang, Shijie Wang, Jun Tang, and 1 others. 2025.
Qwen2. 5-vl technical report.
arXiv preprint arXiv:2502.13923.
Bradley and Terry (1952)
Ralph Allan Bradley and Milton E Terry. 1952.
Rank analysis of incomplete block designs: I. the method of paired comparisons.
Biometrika, 39(3/4):324–345.
Cai et al. (2025)
Yuzheng Cai, Siqi Cai, Yuchen Shi, Zihan Xu, Lichao Chen, Yulei Qin, Xiaoyu Tan, Gang Li, Zongyi Li, Haojia Lin, and 1 others. 2025.
Training-free group relative policy optimization.
arXiv preprint arXiv:2510.08191.
comfyanonymous (2025)
comfyanonymous. 2025.
Comfyui: A modular graph-based interface for diffusion workflows.
https://github.com/comfyanonymous/ComfyUI.
Commit #⟨commit-hash⟩, accessed: YYYY-MM-DD.
Davies et al. (2022)
Simon R Davies, Richard Macfarlane, and William J Buchanan. 2022.
Napierone: A modern mixed file data set alternative to govdocs1.
Forensic Science International: Digital Investigation, 40:301330.
Dvivedi et al. (2024)
Shubhang Shekhar Dvivedi, Vyshnav Vijay, Sai Leela Rahul Pujari, Shoumik Lodh, and Dhruv Kumar. 2024.
A comparative analysis of large language models for code documentation generation.
In Proceedings of the 1st ACM international conference on AI-powered software, pages 65–73.
Fu et al. (2024)
Yuzheng Fu, Yulin Zhang, Yuchen Li, Ruoyu Zheng, and Qiming Liu. 2024.
Aesthetic prediction of mobile app user interface by hybrid deep learning.
IEEE Transactions on Multimedia.
Garfinkel et al. (2009)
Simson Garfinkel, Paul Farrell, Vassil Roussev, and George Dinolt. 2009.
Bringing science to digital forensics with standardized forensic corpora.
digital investigation, 6:S2–S11.
Hong et al. (2024)
Sirui Hong, Mingchen Zhuge, Jonathan Chen, Xiawu Zheng, Yuheng Cheng, Ceyao Zhang, Jinlin Wang, Zili Wang, Steven Ka Shing Yau, Zijuan Lin, and 1 others. 2024.
Metagpt: Meta programming for a multi-agent collaborative framework.
International Conference on Learning Representations, ICLR.
Hui et al. (2024)
Binyuan Hui, Jian Yang, Zeyu Cui, Jiaxi Yang, Dayiheng Liu, Lei Zhang, Tianyu Liu, Jiajun Zhang, Bowen Yu, Keming Lu, Kai Dang, Yang Fan, Yichang Zhang, An Yang, Rui Men, Fei Huang, Bo Zheng, Yibo Miao, Shanghaoran Quan, and 5 others. 2024.
Qwen2.5-coder technical report.
Preprint, arXiv:2409.12186.
Hurst et al. (2024)
Aaron Hurst, Adam Lerer, Adam P Goucher, Adam Perelman, Aditya Ramesh, Aidan Clark, AJ Ostrow, Akila Welihinda, Alan Hayes, Alec Radford, and 1 others. 2024.
Gpt-4o system card.
arXiv preprint arXiv:2410.21276.
Jiang et al. (2024)
Zhouqiang Jiang, Bowen Wang, Junhao Chen, and Yuta Nakashima. 2024.
Relayout: Towards real-world document understanding via layout-enhanced pre-training.
arXiv preprint arXiv:2410.10471.
Li et al. (2018)
Chenglong Li, Jiansheng Wang, Song Yan, Chen Song, and Lianyi He. 2018.
A-lamp: Adaptive layout-aware multi-patch deep cnn for photo aesthetic assessment.
In Proceedings of the 26th ACM international conference on Multimedia, pages 177–185.
Li et al. (2023)
Yuxin Li, Qinglin Zhang, Chen Fang, Jin Xiao, and Wen Xu. 2023.
Towards aligned layout generation via diffusion model with aesthetic constraints.
In The Eleventh International Conference on Learning Representations (ICLR).
Lin et al. (2023)
Jiawei Lin, Jiaqi Guo, Shizhao Sun, Zijiang Yang, Jian-Guang Lou, and Dongmei Zhang. 2023.
Layoutprompter: Awaken the design ability of large language models.
Advances in Neural Information Processing Systems, 36:43852–43879.
Liu and Yu (2023)
Chang Liu and Han Yu. 2023.
Ai-empowered persuasive video generation: A survey.
ACM Computing Surveys, 55(13s):1–31.
Marsili et al. (2025)
Damiano Marsili, Rohun Agrawal, Yisong Yue, and Georgia Gkioxari. 2025.
Visual agentic ai for spatial reasoning with a dynamic api.
In Proceedings of the Computer Vision and Pattern Recognition Conference (CVPR), pages 19446–19455.
Ouyang et al. (2022)
Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, and 1 others. 2022.
Training language models to follow instructions with human feedback.
Advances in neural information processing systems, 35:27730–27744.
Peng et al. (2023)
Sida Peng, Eirini Kalliamvakou, Peter Cihon, and Mert Demirer. 2023.
The impact of ai on developer productivity: Evidence from github copilot.
arXiv preprint arXiv:2302.06590.
Rafailov et al. (2023)
Rafael Rafailov, Archit Sharma, Eric Mitchell, Christopher D Manning, Stefano Ermon, and Chelsea Finn. 2023.
Direct preference optimization: Your language model is secretly a reward model.
Advances in neural information processing systems, 36:53728–53741.
Shao et al. (2024)
Zhihong Shao, Peiyi Wang, Qihao Zhu, Runxin Xu, Junxiao Song, Xiao Bi, Haowei Zhang, Mingchuan Zhang, YK Li, Yang Wu, and 1 others. 2024.
Deepseekmath: Pushing the limits of mathematical reasoning in open language models.
arXiv preprint arXiv:2402.03300.
Stiennon et al. (2020)
Long Ouyang Stiennon, Jeffrey Wu, Daniel M Ziegler, Ryan Lowe, Jeremy Voss, Alec Radford, Dario Amodei, and Paul F Christiano. 2020.
Learning to summarize from human feedback.
Advances in Neural Information Processing Systems, 33:11345–11358.
Subramani et al. (2020)
Nishant Subramani, Alexandre Matton, Malcolm Greaves, and Adrian Lam. 2020.
A survey of deep learning approaches for ocr and document understanding.
arXiv preprint arXiv:2011.13534.
Tang et al. (2023)
Zecheng Tang, Chenfei Wu, Juntao Li, and Nan Duan. 2023.
Layoutnuwa: Revealing the hidden layout expertise of large language models.
arXiv preprint arXiv:2309.09506.
Tian et al. (2025)
Jiaxu Tian, Xuehui Yu, Yaoxing Wang, Pan Wang, Guangqian Guo, and Shan Gao. 2025.
Relayout: Integrating relation reasoning for content-aware layout generation with multi-modal large language models.
arXiv preprint arXiv:2507.05568.
Xie et al. (2024)
Jian Xie, Kai Zhang, Jiangjie Chen, Tinghui Zhu, Renze Lou, Yuandong Tian, Yanghua Xiao, and Yu Su. 2024.
Travelplanner: A benchmark for real-world planning with language agents.
arXiv preprint arXiv:2402.01622.
Xu et al. (2020)
Yiheng Xu, Minghao Li, Lei Cui, Shaohan Huang, Furu Wei, and Ming Zhou. 2020.
Layoutlm: Pre-training of text and layout for document image understanding.
In Proceedings of the 26th ACM SIGKDD international conference on knowledge discovery & data mining, pages 1192–1200.
Yu et al. (2019)
Huimin Yu, Meng Wu, and Bo Cui. 2019.
Calista: A deep learning-based system for understanding and evaluating website aesthetics.
ACM Transactions on Information Systems (TOIS), 37(4):1–27.
Zhang et al. (2024)
Xueru Zhang, Chen Peng, Xin Liao, Shugong Hou, Yang Fan, Li Liang, Yaohui Zhao, and Chunhui Zhang. 2024.
Aesthetiq: Enhancing graphic layout design via aesthetic-aware preference alignment of multi-modal large language models.
In The Twelfth International Conference on Learning Representations (ICLR).
Zheng et al. (2024)
Yaowei Zheng, Richong Zhang, Junhao Zhang, Yanhan Ye, Zheyan Luo, Zhangchi Feng, and Yongqiang Ma. 2024.
Llamafactory: Unified efficient fine-tuning of 100+ language models.
arXiv preprint arXiv:2403.13372.
Zheng et al. (2025)
Ziwei Zheng, Michael Yang, Jack Hong, Chenxiao Zhao, Guohai Xu, Le Yang, Chao Shen, and Xing Yu. 2025.
Deepeyes: Incentivizing" thinking with images" via reinforcement learning.
arXiv preprint arXiv:2505.14362.
Appendix A Appendix
A.1 The Use of Large Language Models
Following the completion of the draft by the human authors, a large language model was employed to enhance the clarity and academic tone of specific sections.
A.2 Model Implementation Details
Our document reward model is built upon the Qwen2.5-VL multimodal architecture, and a regression head is added to predict scalar scores. The maximum input pixels are set to 300,000. It is configured with a maximum context length of 16,000 tokens. Training utilizes the AdamW optimizer with a learning rate of 1e-6 and a batch size of 256 over 3 epochs. All training was conducted on 8 NVIDIA A100 GPUs. The training code is based on LLaMA-Factory (Zheng et al., 2024).
A.3 Source Documents Expansion
To ensure that the reward model learns to assess differences in structure and style rather than content, we applied a rigorous filtering process. Using python-docx, we extracted text from pairs of Microsoft Word DOCX documents and computed their word counts. Only synthetic documents with a word count difference of no more than 20 words from the original document and a ROUGE-L score exceeding a threshold are retained, ensuring comparable content while isolating variation in structure and style. For the constructed training dataset DocPair, both GPT-4o and GPT-5 serve as the base models of agents.
A.4 Annotation Protocol and Reliability
Annotation Guidelines.
The detailed guideline for human annotation are presented in Figure 7.
The annotation guidelines consist of general principles that are formulated in an explicit, objective manner. For instance, extremely narrow margins that produce an almost fully saturated page layout are commonly regarded as unprofessional across different cultural and regional contexts.
Figure 7: Detailed guideline for human annotation.
Independence from annotators’ cultural and professional backgrounds.
The annotation was performed by three Ph.D. students with complementary expertise in computer science/math, marketing, and design.
We measured inter-annotator reliability using Cohen’s Kappa; the results are shown in Table 7.
The high agreement indicates that the annotations follow clear, well-defined rules that do not depend on the annotators’ professional training or cultural background, demonstrating the guidelines’ generality and objectivity.
Annotator 1
Annotator 2
Annotator 3
Average
Annotator 1
83.40
80.92
82.15
Annotator 2
83.40
85.90
84.65
Annotator 3
80.92
85.90
83.41
Average
82.15
84.65
83.41
83.40
Table 7: Cohen’s Kappa among annotators.
A.5 Document Types Distribution
The top 30 document types are presented in Figure 8.
Figure 8: Top 30 Document Type Distribution.
A.6 Prompts
A.7 Details of Best-of-N Evaluation
For the Best-of-N experiment, the Textual Content to Document defined in Section 3 is adopted as the document agent, with the base model being GPT-5. Three reward models, including random, GPT-5, and DocReward are compared. Once the document agent generates candidates and the reward model selects the top-ranking document from candidates, a highly educated annotator is asked to rank the three documents selected, according to the definitions of professional structure and style defined in Figure 7. As a result, documents from each reward model are annotated 130 comparison pairs against those of another reward model.
Finally, the win/lose/tie rate of each reward model is calculated on the comparison pairs against the other reward models.
A.8 Ablation Study of Inputs
In designing the input channels for DocReward, we experimented with two different configurations: a purely visual channel method and a combination method of visual and additional parsing information. The experimental results are summarized in Table 8. Results show that text and bounding box of text span from an additional OCR module are not helpful for the assessment of professional structure and style.
Inputs
Accuracy
image-only (3B)
85.00
image + OCR text & bbox (3B)
80.30
image-only (7B)
87.94
image + OCR text & bbox (7B)
84.41
Table 8: Additional text and bounding box of text span are not helpful for the assessment of professional structure and style.
A.9 Case Study
We present a case study on documents with identical textual content but different structures and styles in Figure 9. In case (a), the allocation of whitespace is ineffective, with insufficient space between Last Name and excessive space between First Name, leading to an imbalanced layout. Key fields such as Faculty/Department, Country, and Country Code are not vertically aligned, causing a cluttered and disorganized layout. This poor alignment and inconsistent spacing result in a low score of 1.21 from DocReward. Case (b) adopts a table-like arrangement, but the level-1 heading The teaching staff member is too small and does not stand out from the body text, diminishing its impact. Additionally, the lack of borders around input fields makes it hard to locate items easily, resulting in a moderate score of 2.11. Case (c) provides a clear and well-structured layout, with headings appropriately larger than the body text and better readability, earning the highest rating of 5.34. These results show that DocReward effectively captures document professionalism in structure and style.
(a) score: 1.21
(b) score: 2.11
(c) score: 5.34
Figure 9: Case study: DocReward’s scores reflect structural and stylistic professionalism.